0.5 kg of air is taken through a constant pressure cycle. At the beginning of
adiabatic compression are 96.5 kN/m 2 and 15 0 C. The pressure ratio of
compression is 6. Constant pressure heat addition occurs after adiabatic
compression until the volume is doubled. Take γ = 1.4 and R= 0.289 kJ/kg K.
Determine for the cycle:
(a) The pressure, volume and temperature at cycle state points. (b) The heat received.
(c) The thermal efficiency. (d) The work ratio.
(e) The heat rejected.
1
Expert's answer
2020-09-16T05:13:50-0400

mass of air =0.5 kg, P1=96.5 kN/m2,T1= 15 0C,compression ratio=6, Heat addition takes place at constant pressure to volume become doubled.

(a)

"\frac{P_2}{P_1}= r_p" ,

"P_2= r_p\times P_1=6\times 96.5= 579 \frac{kN}{m^2}"

"\frac{T_1}{T_2}=(\frac{P_1}{P_2})^{\frac{\gamma -1}{\gamma}}"


"\frac{288}{T_2}=(\frac{96.5}{579})^{\frac{1.4 -1}{1.4}}"


"T_2=480.8 K"

"\frac{P_1}{P_2}^{\frac{1}{\gamma}}= \frac{V_2}{V_1}"

for finding v1 we use ideal gas equation

"P_1V_1=nRT_1"

"95.5\times V_1=1\times 0.289\times 288"

"V_1=0.8625 m^3"


"\frac{96.5}{579}^{\frac{1}{1.4}}= \frac{V_2}{0.8625}"


"V_2=0.2398 m^3"


(b)"H_1= m C_p \Delta T=0.5\times 1.0115\times 192.8= 97.508 kJ"

(c) Thermal efficiency

"\eta_t= 1- (\frac{V_2}{V_1})^{\gamma -1}"


"\eta = 0.40"

(d)

Work ratio="1- (\frac{T_1}{T_2})\times r^{\frac{\gamma -1}{\gamma}}"


Work ratio= 0.01

for heat rejected

"\eta = 1- \frac{Q_2}{Q_1}"


"Q_2= 58.5 kJ" it is heat rejected